1) Ben Bitdiddle is building a house. Explain how he can use the principles of hierarchy, modularity, and regularity to save time and money during construction. 2) Convert the following two’s complement binary numbers to decimal. (a) 1010 (b) 110110 (c) 01110000 (d) 10011111
Answer #1

2)
converting binary numbers to decimal:
a)
1010 = 1*2^3 +0*2^2 +1*2^1 +0*2^0 = 8+2 = 10 in decimal
b)
110110 = 1*2^5 +1*2^4 +0*2^3 +1*2^2 +1*2^1 +0*2^0 = 32+16+4+2 = 54
c)
01110000 = 0*2^7 +1*2^6 +1*2^5 +1*2^4 +0*2^3 +0*2^2 +0*2^1 +0*2^0 = 64+32+16 =112
d)
10011111 = 1*2^7 +0*2^6 +0*2^5 +1*2^4 +1*2^3 +1*2^2 +1*2^1 +1*2^0 = 128+16+8+4+2+1=159

Note: As per Chegg policy, I am allowed to answer only 1 question(including sub-parts)
on a single post, kingly post the remaining questions seperately and i will try to answer them
Sorry for the inconvenience caused, thank you